OpenStudy (anonymous):

Write an equation in slope-intercept, point-slope, or standard form for the line with the given information. Explain why you chose the form you used. a. passes through (-1, 4) and (-5, 2 b. slope 2, y-intercept -4 c. has an x-intercept of 6 and a y-intercept of 3 d. passes through (1, 2) with slope -5/3

OpenStudy (anonymous):

well we have different forms of equation of a st.line : =============================== >> if you are given two points then it will be in the form \[\frac{ y-y1 }{ y2-y1 } = \frac{ x-x1 }{x2-x1 }\] where the two points are (x1,y1) ^ (x2,y2) >> if you are given the slope and the y-intercept \[y=mx+c\] where m is the slope and c is the y-intercept >> if you are given a point and the slope \[y-y1=m(x-x1)\] >> AM NOT SURE OF THIS I guess if you are given y-intercept and x-intercept so implicitly we are given two points we can consider it like the case of two given points I recommend you check the last one cuz am not certain about it